OpenStudy (anonymous):

The ages of A and B are of the ratio 2:5. 5 years ago they were in the ratio of 1:10. Find the age of B?

OpenStudy (raden):

A/B = 2/5 ---> A = 2B/5 ... (1) (A-5)/(B-5) = 1/10 ... (2) subtitute the 1st equation into 2nd equation then solve for B

OpenStudy (anonymous):

dont get it

OpenStudy (raden):

(A-5)/(B-5) = 1/10 now change A = 2B/5 to get (2B/5 - 5)/(B - 5) = 1/10 cross multiplication! 10 * (2B/5 - 5) = B - 5 10 * 2B/5 - 10 * 5 = B - 5 4B - 50 = B - 5 now can solve for B ?

OpenStudy (raden):

*can you

OpenStudy (anonymous):

i dont understand the way you have written it.

OpenStudy (raden):

which parts you still dont get it ?

OpenStudy (loser66):

can I?

OpenStudy (anonymous):

yep

OpenStudy (raden):

sure :)

OpenStudy (loser66):

let say 5years ago, age of A is x age of B is 10x does it make sense?

OpenStudy (loser66):



OpenStudy (loser66):

now, it means after 5 years, age of A is x +5 age of B is 10x +5 got this part?

OpenStudy (loser66):